# 4-LEGS: 4D Language Embedded Gaussian Splatting

### v2

@@ 2 Related Work
Several recent works embed language onto 3D Gaussians representations [QLZ23, ZCJ23, SWDG23], which achieve real-time inference, significantly outperforming NeRF-based techniques. As inference speed is critical for enabling an interactive interface for spatiotemporally localizing events in videos, in our work, we also embed language onto Gaussians. However, unlike prior work, we focus on the modeling of dynamic scenes, augmenting a dynamic 3D Gaussian representation [LKLR23] with high-dimensional language features. Additionally, several work [KWK24, YDYK23, CFY23, HWF24] segment 3D scenes via high-dimensional features optimized by leveraging 2D masks provided, for instance, by Segment Anything [KMR23]. However, these works do not directly utilize language, nor do they allow for text-based querying.
@@ 4.1 The Grounding-PanopticSports Benchmark
Metrics. We adopt a set of evaluation metrics that consider both the spatial and temporal dimensions, as well as the 3D nature of our multi-view data. For evaluating bounding box outputs we follow 2D techniques (e.g.  [ZZZ20b]) and define $vIoU$ as
@@
To separately evaluate the quality of the temporal localization, we also measure Recall (tRec), Precision (tPrec), and temporal IoU (tIoU) for each video, and average these across all annotated videos. Additionally, to provide a measure for consistency across views, we also report the standard deviation among the different views over all metrics, averaging over the different scenes.
@@ 4.2 Baselines
For the 2D spatio-temporal grounding baseline methods, we consider TubeDeTR [YMS22] and Context-Guided Spatio-Temporal Video Grounding (CGSTVG) [GFH24]. TubeDETR uses visual and text encoders to create video-text features based on short clips extracted from the video. A space-time decoder is then used to predict both bounding boxes for each frame and the start and end frame by using temporal self-attention to model the long-range temporal interactions in the video and masked cross-attention. CGSTVG uses a multimodal encoder for feature extraction and a context-guided decoder by cascading a set of decoding stages for grounding. In each decoding stage, instance context is mined to guide query learning for better localization. The output of these methods is a sequence of bounding boxes matching to each frame of the estimated temporal tube. As these 2D methods predict bounding boxes (and not segmentations), we convert them to segmentation maps by filling in ones inside each box and zeros outside of it.
For the comparison to a 3D language embedding technique, we compare with LangSplat [QLZ23], which optimizes a per-scene 3D static language field on a 3D Gaussian Splatting [KKLD23] representation using hierarchical SAM [KMR23] masks and CLIP [RKH21] features. To adapt this method for our temporal scenes, we optimize a LangSplat per timestep (as LangSplat only operates over static scenes). We optimize LangSplats for a small number of timesteps, as exhaustively optimizing LangSplat models across all timestamps and scenes would be computationally intractable. As one LangSplat is made up of three 3DGS language fields, each for a different semantic scale, we average the relevancy map computed from each scale to create a single output for comparison. For the evaluation of LangSplat over $vIoU$, we convert the predicted segmentation maps into bounding boxes by taking the connected component with the maximum score (we perform this for our method as well).
@@ 4.3 Comparisons
A quantitative evaluation on the Grounding-PanopticSports benchmark is provided in Table 1. As illustrated in the table (and also qualitatively in Figure 2), our method outperforms the baseline 2D methods significantly, yielding better results across all metrics. In particular, their results highly vary (high standard deviation value). In other words, the predictions of 2D methods are inconsistent across different views. This is further illustrated in Figure 2 on the right, where different predictions are extracted for differing views during the same timestep. Our method achieves consistency between different views as it optimizes a volumetric representation.
Figure 5 shows a qualitative evaluation to LangSplat. As illustrated in Table 1 and in the figure, our method outperforms LangSplat, emphasizing the importance of distilling temporal features in addition to spatial features. Figure 5 shows that LangSplat cannot localize the input queries temporally, and that it outputs high probabilities for all timesteps. It also demonstrates that LangSplat yields high predicted probabilities even for input images do not depict the queried text. This is further observed in our ablations, where we can see that CLIP-based features struggle in localizing actions and instead focus on objects. Our method, by contrast, is capable of temporally localizing an input query, due to the use of temporal features, as illustrated both numerically and in the figure. In the supplementary, we show results over additional views.
@@
TubeDETR $27.6\pm 5.2$ $19.9\pm 7.3$ $29.3\pm 9.7$ $44.6\pm 16.4$ $63.7\pm 16.3$
CGSTVG $32.8\pm 10.1$ $24.3\pm 7.7$ $34.9\pm 10.8$ $56.7\pm 17.3$ $51.6\pm 11.3$
LangSplat $35.8\pm 0.3$ $18.2\pm 3.9$
Ours $\mathbf{58.7\pm 0.6}$ $\mathbf{27.0\pm 4.2}$ $\mathbf{60.8\pm 0.0}$ $\mathbf{83.0\pm 0.0}$ $\mathbf{68.5\pm 0.0}$
Table 1: Quantitative Evaluation over the Grounding-PanopticSports Benchmark. We report performance over the spatio-temporal vAP metric, and over three additional metrics quantifying the quality of the temporal localization (tIOU,tRec, tPrec). Note that LangSplat metrics are calculated over a small number of timesteps, as further detailed in Section 4.2.
@@ 4.4 Ablations
We ablate our use of the ViCLIP video encoder, demonstrating the importance of extracting features with a video encoder, our use of a volumetric representation, emphasizing the benefits of lifting features from 2D. When ablating our use of a video encoder we compare against two different methods for extracting pixel-aligned 2D spatio-temporal features: (i) static CLIP features (denoted as $\text{Static}_{\text{CLIP}}$), which uses a CLIP image encoder and does not look at a temporal tube centered at the current frame, and (ii) averaged CLIP features (denoted as $\text{Avg}_{\text{CLIP}}$), which averages the extracted CLIP image encoder features in a temporal window centered at our current frame. When ablating the use of a volumetric representation we compare against our extracted pixel-aligned 2D features (denoted as 2D Features). As illustrated in Figure 4, CLIP-based methods struggle to localize the predictions temporally, even when using averaged features. Quantitatively, as illustrated in Table 2, our method achieves significantly higher scores across all metrics. The 2D Features baseline is also outperformed by our method. Additionally, results highly vary across different viewpoints. This is in contrast to consistency achieved by our method, due to the use a volumetric representation. In the supplementary material, we conduct additional experiments, ablating our feature extraction aggregation and our attention mechanism. These ablations demonstrate that our averaging mechanism outperforms other aggregation techniques and that our attention mechanism improves viewpoint consistency and produces smoother predicitions.
@@ 4.5 Limitations
Our method can localize a given text query spatially and temporally. However, there are several limitations to consider. First, as the 3D Gaussian Splatting representation we build upon is optimized per timestep, runtime may potentially become an issue for longer videos. We also inherit its quality for rendering the scene from novel views. In the spatio-temporal highlights we created, we did observe some visual artifacts, particularly when the rendered Gaussians are close to the camera view. Second, our localization accuracy is tied to our video encoder backbone. We found that ViCLIP struggled with fine grained localization of some actions (e.g., highlighting the entire Panda rather than its paws in Figure 7). Given the high volume of research focusing on video-to-text models as well as field-based representations, we expect both localization accuracy and video quality to improve with future backbones.
@@
[…]
2D Features $44.9\pm 11.7$ $46.3\pm 12.0$ $73.3\pm 14.2$ $53.7\pm 11.8$
$\text{Static}_{\text{CLIP}}$ $33.3\pm 1.0$ $34.9\pm 0.0$ $62.2\pm 0.0$ $43.4\pm 00.0$
$\text{AVG}_{\text{CLIP}}$ $32.5\pm 0.7$ $35.2\pm 0.0$ $62.5\pm 0.0$ $43.6\pm 0.0$
Ours $\mathbf{58.7\pm 0.6}$ $\mathbf{60.8\pm 0.0}$ $\mathbf{83.0\pm 0.0}$ $\mathbf{68.5\pm 0.0}$
Table 2: Ablation study, ablating the use of a 4D learned feature field (2D Features) and the video–text model ($\text{Static}_{\text{CLIP}}$, $\text{AVG}_{\text{CLIP}}$). Best results are highlighted in bold.

### v3

@@ 2 Related Work
Several recent works embed spatial language features onto 3D Gaussians representations [QLZ23, ZCJ23, SWDG23, LILB24], which achieve real-time inference, significantly outperforming NeRF-based techniques. As inference speed is critical for enabling an interactive interface for spatiotemporally localizing events in videos, in our work, we also embed language onto Gaussians. However, unlike these prior works, we focus on the modeling of dynamic scenes, augmenting a dynamic 3D Gaussian representation [LKLR23] with high-dimensional spatio-temporal language features. Additionally, several work [KWK24, YDYK23, CFY23, HWF24] segment 3D scenes via high-dimensional features optimized by leveraging 2D masks provided, for instance, by Segment Anything [KMR23]. However, these works do not directly utilize language, nor do they allow for text-based querying.
@@ 4.1 The Grounding-PanopticSports Benchmark
Metrics. We adopt a set of evaluation metrics that consider both the spatial and temporal dimensions, as well as the 3D nature of our multi-view data. We follow 2D techniques (e.g.  [ZZZ20b]) and define $vIoU$ as
@@
To separately evaluate the quality of the temporal localization, we also measure Recall (tRec), Precision (tPrec), AP (tAP) and temporal IoU (tIoU) for each video, and average these across all annotated videos. Additionally, to provide a measure for consistency across views, we also report the standard deviation among the different views over all metrics, averaging over the different scenes.
@@ 4.2 Baselines
For the 2D spatio-temporal grounding baseline methods, we consider TubeDeTR [YMS22] and Context-Guided Spatio-Temporal Video Grounding (CGSTVG) [GFH24]. TubeDETR uses visual and text encoders to create video-text features based on short clips extracted from the video. A space-time decoder is then used to predict both bounding boxes for each frame and the start and end frame by using temporal self-attention to model the long-range temporal interactions in the video and masked cross-attention. CGSTVG uses a multimodal encoder for feature extraction and a context-guided decoder by cascading a set of decoding stages for grounding. In each decoding stage, instance context is mined to guide query learning for better localization. The output of these methods is a sequence of bounding boxes matching to each frame of the estimated temporal tube.
For the comparison to a 3D language embedding technique, we compare with LangSplat [QLZ23], which optimizes a per-scene 3D static language field on a 3D Gaussian Splatting [KKLD23] representation using hierarchical SAM [KMR23] masks and CLIP [RKH21] features. To adapt this method for our temporal scenes, we optimize a LangSplat per timestep (as LangSplat only operates over static scenes). We optimize LangSplats for a small number of timesteps, as exhaustively optimizing LangSplat models across all timestamps and scenes would be computationally intractable. As one LangSplat is made up of three 3DGS language fields, each for a different semantic scale, we average the relevancy map computed from each scale to create a single output for comparison.
@@ 4.3 Comparisons
Temporally localizing actions presents unique challenges as actions span multiple timesteps and can occur multiple times. Our 4D representation addresses these challenges in ways that 2D and 3D baselines cannot.
A quantitative evaluation on the Grounding-PanopticSports benchmark is provided in Table 1. As illustrated in the table (and also qualitatively in Figure 2), our method outperforms the baseline 2D methods significantly, yielding better results across all metrics. The 2D methods suffer from view inconsistency as they infer over a single view, failing to utilize information from other viewpoints. In particular, their results highly vary (high standard deviation value). This is further illustrated in Figure 2 on the right, where different predictions are extracted for differing views during the same timestep. Our method achieves consistency between different views as it optimizes a volumetric representation in 4D space, thus utilizing data from all views.
Figure 5 shows a qualitative evaluation to LangSplat. As illustrated in Table 1 and in the figure, our method outperforms LangSplat, emphasizing the importance of distilling temporal features in addition to spatial features. LangSplat suffers from lack of temporal awareness as it is optimized independently for each timestep. Figure 5 shows that LangSplat cannot localize the input queries temporally, and that it outputs high probabilities for all timesteps. It also demonstrates that LangSplat yields high predicted probabilities even for input images do not depict the queried text. 4-LEGS achieves temporal awareness and proper temporal localization of actions, by optimizing over all timesteps.
This is further observed in our ablations, where we can see that CLIP-based features struggle in localizing actions and instead focus on objects. Our method, by contrast, is capable of temporally localizing an input query, due to the use of temporal features, as illustrated both numerically and in the figure. In the supplementary, we show results over additional views.
@@
TubeDETR $21.3\pm 3.7$ $19.9\pm 7.3$ $29.3\pm 9.7$ $44.6\pm 16.4$ $63.7\pm 16.3$ $41.4\pm 13.7$
CGSTVG $25.1\pm 8.3$ $24.3\pm 7.7$ $34.9\pm 10.8$ $56.7\pm 17.3$ $51.6\pm 11.3$ $48.1\pm 14.4$
LangSplat $38.4\pm 0.3$ $13.9\pm 4.1$
Ours $\mathbf{58.7\pm 0.6}$ $\mathbf{25.6\pm 4.6}$ $\mathbf{60.8\pm 0.0}$ $\mathbf{83.0\pm 0.0}$ $\mathbf{68.5\pm 0.0}$ $\mathbf{72.6\pm 0.0}$
Table 1: Quantitative Evaluation over the Grounding-PanopticSports Benchmark. We report performance over the spatio-temporal vAP and vIOU metrics, and over four additional metrics quantifying the quality of the temporal localization (tIOU,tRec, tPrec, tAP). Note that LangSplat metrics are calculated over a small number of timesteps, as further detailed in Section 4.2.
@@ 4.4 Ablations
We ablate our use of the ViCLIP video encoder, demonstrating the importance of extracting features with a video encoder, our use of a volumetric representation, emphasizing the benefits of lifting features from 2D. When ablating our use of a video encoder we compare against two different methods for extracting pixel-aligned 2D spatio-temporal features: (i) static CLIP features (denoted as $\text{Static}_{\text{CLIP}}$), which uses a CLIP image encoder and does not look at a temporal tube centered at the current frame, and (ii) averaged CLIP features (denoted as $\text{Avg}_{\text{CLIP}}$), which averages the extracted CLIP image encoder features in a temporal window centered at our current frame. When ablating the use of a volumetric representation we compare against the extracted ViCLIP features from the input 2D videos (denoted as 2D Features). The CLIP-based methods suffer from poor temporal localization since they distill only static features. This is illustrated in Figure 4, as these methods struggle to localize the predictions temporally, even when using averaged features. In contrast 4-LEGS distills temporally-aware features thus enabling correct temporal localization. This is further illustrated quantitatively in Table 2, as our method achieves significantly higher scores across all metrics. The 2D Features baseline (the extracted ViCLIP features from the input 2D videos) is also outperformed by our method. Additionally, results highly vary across different viewpoints. This is in contrast to consistency achieved by our method, due to the use a volumetric representation.
Additionally, we ablate the use of self-attention and its effectiveness in enforcing local smoothness in 3D space between the disttilled features. We compare our use of self-attention against a 4D language feature field distilled without attention (denoted as w/o Attn) and its effectiveness in enforcing smoothness within a local spatial neighborhood with three methods of local smoothing, a neighborhood blur (denoted as $\text{Local}_{\text{Blur}}$), a neighborhood smoothness regularization loss (denoted as $\text{Local}_{\text{Reg}}$) and a MLP for the local neighborhood (denoted as $\text{Local}_{\text{MLP}}$. As illustrated in Table 3, the addition of self-attention improves performance and it’s ability to model local relationships is outperforms all other local smoothness operators. We provide additional details and a qualitative comparison in the supplementary material.
In the supplementary material, we conduct additional experiments, ablating our feature extraction aggregation , our temporal localization score weighting method and the querying space usef for open-vocabulary querying. These ablations demonstrate that our averaging mechanism outperforms other aggregation techniques , that weighting the temporal localization scores by cardinality outperforms weighting by size and that querying in the feature space outperforms latent space querying.
@@ 4.5 Limitations
Our method can localize a given text query spatially and temporally. However, there are several limitations to consider. First, as the 3D Gaussian Splatting representation we build upon is optimized per timestep, runtime may potentially become an issue for longer videos. We also inherit its quality for rendering the scene from novel views. In the spatio-temporal highlights we created, we did observe some visual artifacts, particularly when the rendered Gaussians are close to the camera view. Second, our localization accuracy is tied to our video encoder backbone. We found that ViCLIP struggled with fine grained localization of some actions (e.g., highlighting the entire Panda rather than its paws in Figure 7). Given the high volume of research focusing on video-to-text models as well as field-based representations, we expect both localization accuracy and video quality to improve with future backbones. Third, training both dynamic 3D Gaussians and an autoencoder per scene is time-consuming and computationally intense. Speeding up preprocessing is promising future work.
[…]
2D Features $44.9\pm 11.7$ $18.6\pm 10.3$ $46.3\pm 12.0$ $73.3\pm 14.2$ $53.7\pm 11.8$ $60.3\pm 12.5$
$\text{Static}_{\text{CLIP}}$ $33.3\pm 1.0$ $8.2\pm 3.9$ $34.9\pm 0.0$ $62.2\pm 0.0$ $43.4\pm 00.0$ $47.2\pm 0.0$
$\text{AVG}_{\text{CLIP}}$ $32.5\pm 0.7$ $7.4\pm 3.7$ $35.2\pm 0.0$ $62.5\pm 0.0$ $43.6\pm 0.0$ $48.4\pm 0.0$
Ours $\mathbf{58.7\pm 0.6}$ $\mathbf{25.6\pm 4.6}$ $\mathbf{60.8\pm 0.0}$ $\mathbf{83.0\pm 0.0}$ $\mathbf{68.5\pm 0.0}$ $\mathbf{72.6\pm 0.0}$
Table 2: Ablation study, ablating the use of a 4D learned feature field (2D Features, the extracted ViCLIP features from the input 2D videos) and the video–text model ($\text{Static}_{\text{CLIP}}$, $\text{AVG}_{\text{CLIP}}$). Best results are highlighted in bold.
[…]
$\text{Local}_{\text{Blur}}$ $55.3\pm 1.3$ $19.8\pm 5.1$ $57.9\pm 0.0$ $83.1\pm 0.0$ $65.2\pm 0.0$ $70.3\pm 0.0$
$\text{Local}_{\text{Reg}}$ $56.7\pm 1.4$ $21.1\pm 4.7$ $59.3\pm 0.0$ $82.7\pm 0.0$ $66.4\pm 0.0$ $71.1\pm 0.0$
$\text{Local}_{\text{MLP}}$ $51.7\pm 1.2$ $18.2\pm 3.3$ $54.1\pm 0.0$ $\mathbf{85.4\pm 0.0}$ $59.6\pm 0.0$ $67.7\pm 0.0$
w/o Attn $57.9\pm 1.6$ $20.1\pm 4.2$ $60.6\pm 0.0$ $83.6\pm 0.0$ $67.7\pm 0.0$ $67.9\pm 0.0$
Ours $\mathbf{58.7\pm 0.6}$ $\mathbf{25.6\pm 4.6}$ $\mathbf{60.8\pm 0.0}$ $83.0\pm 0.0$ $\mathbf{68.5\pm 0.0}$ $\mathbf{72.6\pm 0.0}$
Table 3: Self-Attention Ablations, ablating the use of self-attention (w/o Attn) and the effectiveness of self-attention in enforcing local smoothness in the 3D space ($\text{Local}_{\text{Blur}}$, $\text{Local}_{\text{Reg}}$ and $\text{Local}_{\text{MLP}}$). Best results are highlighted in bold. See the supplementary material for additional details.
@@ 5 Applications
Input query: A person stretching
Input query: A person bending
Figure 9: Qualitative results on the Neural Human Rendering [WWHY20] dataset, with the input frames depicted on top and the rendered probabilities directly below (frames localized temporally are in red). As illustrated above, our approach succeeds in temporally localizing human actions. We show results over additional scenes in the supplementary material.